# The NANOGrav 15 year Data Set: Removing pulsars one by one from the pulsar timing array

### v1

@@ Appendix A Removing the noisiest pulsars first
In the main body of this work we show analysis results as pulsars are removed from the array, with the least-noisy pulsars (those with the smallest white noise TOA uncertainties) removed first. This appendix shows the results obtained with the opposite order i.e., when the noisiest pulsars are removed first. Figure 7 shows how the SCNMOS and associated S/N evolve as pulsars are removed, while Figure 8 shows the same for the MCNMOS. In both cases $\hat{A}^{2}$ remains relatively steady until approximately 40 pulsars have been omitted. The S/N is relatively steady until approximately 30 pulsars have been removed.
Figure 7: The evolution of the SCNMOS and S/N for NG15 as the most-noisy pulsars are removed one by one from the PTA. Similar to Figure 2; see that caption for more details. In this case we do not show a shaded region beyond 50 pulsars removed (see e.g., Figure 2). This is because low-noise pulsars remain in the array beyond this point and so the behavior of $\hat{A}^{2}$ and S/N is less erratic.
Figure 8: The evolution of the MCNMOS and S/N for NG15 as the most-noisy pulsars are removed one by one from the PTA. Similar to Figure 2; see that caption for more details. In this case we do not show a shaded region beyond 50 pulsars removed (see e.g., Figure 2). This is because low-noise pulsars remain in the array beyond this point and so the behavior of $\hat{A}^{2}$ and S/N is less erratic.
@@ Appendix B Distribution of Simulated Features
Figure 9: The distribution of $\hat{A}^{2}$ and S/N with all 67 NANOGrav pulsars in the array (Metric 1 from Section 5.1). The histograms shows the median values from the MCNMOS analysis for each of the 100 simulated data sets. The results for monopole, dipole and HD spatial correlations are red, green and blue respectively. The dashed vertical lines indicate the median values from the MCNMOS analysis of NG15.

### v2

@@ Appendix A Removing the noisiest pulsars first
In the main body of this work we show analysis results as pulsars are removed from the array, with the least noisy pulsars (those with the smallest white noise TOA uncertainties) removed first. This appendix shows the results obtained with the opposite order i.e., when the noisiest pulsars are removed first. Figure 7 shows how the SCNMOS and associated S/N evolve as pulsars are removed, while Figure 8 shows the same for the MCNMOS. In both cases $\hat{A}^{2}$ remains relatively steady until approximately 40 pulsars have been omitted. The S/N is relatively steady until approximately 30 pulsars have been removed.
Figure 7: The evolution of the SCNMOS and S/N for NG15 as the most noisy pulsars are removed one by one from the PTA. Similar to Figure 2; see that caption for more details. In this case we do not show a shaded region beyond 50 pulsars removed (see e.g., Figure 2). This is because low-noise pulsars remain in the array beyond this point and so the behavior of $\hat{A}^{2}$ and S/N is less erratic.
Figure 8: The evolution of the MCNMOS and S/N for NG15 as the most noisy pulsars are removed one by one from the PTA. Similar to Figure 2; see that caption for more details. In this case we do not show a shaded region beyond 50 pulsars removed (see e.g., Figure 2). This is because low-noise pulsars remain in the array beyond this point and so the behavior of $\hat{A}^{2}$ and S/N is less erratic.
@@ Appendix B Distribution of Simulated Features
Figure 9: The distribution of $\hat{A}^{2}$ and S/N with all 67 NANOGrav pulsars in the array (Metric 1 from Section 5.1). The histograms shows the median values from the MCNMOS analysis for each of the 100 simulated data sets. The results for the monopole, dipole, and HD spatial correlations are red, green and blue respectively. The dashed vertical lines indicate the median values from the MCNMOS analysis of NG15.